# Two liquid states of distinguishable helium-4: the existence of another non-superfluid frozen by heating

### v2

@@ Abstract
We demonstrate that there can exist two liquid states in distinguishable helium-4 (4He) obeying Boltzmann statistics. This is an indication of quantum liquid polyamorphism induced by nuclear quantum effect. For 0.08-3.3 K and 1-500 bar, we extensively conducted the isothermal-isobaric path integral centroid molecular dynamics simulations to explore not only possible states and state diagram but the state characteristics. The distinguishable 4He below 25 bar does not freeze down to 0.08 K even though it includes no Bosonic exchange effect and therefore no Bose condensation. One liquid state, low quantum-dispersion liquid (LQDL), is nearly identical to normal liquid He-I of real 4He. The other is high quantum-dispersion liquid (HQDL) consisting of atoms with longer quantum wavelength. This is another non-superfluid existing below 0.5 K or the temperatures of LQDL. The HQDL is also a low-entropy and fragile liquid to exhibit, unlike conventional liquids, rather gas-like relaxation of velocity autocorrelation function, while there the atoms diffuse without noticeable contribution from quantum tunneling. The LQDL-HQDL transition is not a thermodynamic phase transition but a continuous crossover accompanied by the change of the expansion factor of quantum wavelength. Freezing of HQDL into the low quantum-dispersion amorphous solid occurs by heating from 0.2 to 0.3 K at 40-50 bar, while this $P$-$T$ condition coincides with the Kim-Chan normal-supersolid phase boundary of real 4He. It is suggested that HQDL has relevance to the non-superfluid states of confined subnano-scale 4He systems which are semi-Boltzmann-like owing to the suppression of Bosonic correlation.
@@ I INTRODUCTION
Is it a true postulation that every substance should freeze into its solid phase as it is cooled down toward absolute zero temperature? While classical systems inevitably freeze by losing kinetic energy proportional to temperature, anti-freezing of quantum systems is benefited by promotion of atomic diffusivity due to increasing quantum atomic fluctuation down to zero temperature. Indeed, we know an exceptionally non-freezing quantum substance, helium-4 (4He), which retains the superfluid (He-II) phase which never freezes down to zero temperature under atmospheric pressure. The cause of superfluidity includes: (i) the nuclear quantum effect (NQE), i.e., the atomic wave nature parametrized as the de Broglie thermal wavelength $\lambda_{\rm{dB}}:=\hbar\sqrt{2\pi/mk_{\rm{B}}T}$ ($m$ is the atomic mass and $k_{\rm{B}}$ is the Boltzmann constant), (ii) the weakness of the van der Waals interatomic forces, and (iii) the atomic indistinguishableness due to the Bosonic permutation [1]. Primitively, the $\rm{\lambda}$-type isobaric heat capacity $C_{P}$ at the transition between normal liquid (He-I) and He-II of 4He is explained in terms of the Bose-Einstein condensation (BEC) of ideal Bose gas [2, 3, 4]. This leads to a conventional understanding that the Bose statistics is the key to the $\lambda$ transition and causes the BEC through which non-freezing of this quantum substance is achieved. Then, if there were solely the NQE and no Bosonic exchange were imposed on atoms, could such a distinguishable 4He model avoid freezing to retain the normal liquid phase when cooled? In Huang’s famous textbook, the non-freezing property of 4He was explained by considering only the two factors (i) the NQE of individual atoms and (ii) the weak interatomic forces [5]. However, Boninsegni et al. recently claimed that the third factor, (iii) the Bosonic exchange among $N$ atoms, was indispensable to retaining the non-freezing property by being in superfluid state and the crystal of distinguishable 4He did not melt at the same temperature-density condition [6]. Therefore, it is not yet completely resolved whether the sole NQE of individual atoms really causes the non-freezing property of quantum substance.
@@
The non-freezing tendency of a quantum substance might be reflected to the ease of residual diffusion [7] of atoms in its quenched glassy solid. In this connection, Markland et al. conducted the discretized path integral simulations based on the quantum-classical isomorphism, where quantum atoms are expressed as classical ring polymers (necklaces), to show the reentrant behavior of self-diffusion coefficient $D$ in the glassy solid of a binary Lennard-Jones model. In their study, increasing the quantumness parameter $\Lambda^{*}:=\hbar/\sqrt{mk_{\rm{B}}T\sigma^{2}}$ ($\sigma$ is the closest interatomic distance) caused the transition from a glass state involving trapped configurations of atomic necklaces (called trapped regime) to another state with stretched configurations of atoms (tunneling regime[8, 9]. As $\Lambda^{*}$ increased, $D$ once decreased in the former regime but it turned to increase in the latter [8, 9]. This also suggested a tendency toward inverse melting (or melting by cooling) of this quantum system [9]. In fact, such an atomic tunneling between local potential minima explains the linear temperature dependence of heat capacity in low-temperature glasses [10, 11].
@@
The existence of more than one states in liquids or glasses is called polyamorphism [12, 13], which is, for example, observed for classical fragile liquids [14] such as water [15, 16] and phosphorous [17]. The existence of two liquid states, He-I and He-II, in 4He is certainly an example of liquid polyamorphism. However, the polyamorphism of the other quantum systems has rarely been discovered so far [18, 19, 20]. Recently, our path integral centroid molcular dynamics (CMD) simulations showed a quantum polyamorphism of compressed distinguishable 4He glass above 3 K [21]. In good accordance with the two regimes reported by Markland et al. [8, 9], compressed distinguishable 4He also involved two glass states [21]. The two states were distinguished by the difference in atomic quantum wavelength $\lambda_{\rm{quantum}}$ which is equivalent to two times radius of gyration of atomic necklaces [21]. Then we call the two states, depending on $\lambda_{\rm{quantum}}$, the low quantum dispersion amorphous solid (LQDA) and the high quantum dispersion solid (HQDA) [21]  [^1] Here we should mention the abbreviation of these names. In Ref.  21, we abbreviated them as “LDA” and “HDA”, respectively. However, since these may be confused with the low and high density amorphous state of classical systems (e.g., water [15, 16]), we presently call them LQDA and HQDA instead.. The LQDA and HQDA are equivalent to the trapped and the tunneling regimes in the quantum Lennard-Jones model [8, 9], respectively. The LQDA-HQDA transition was reversibly caused by varying temperature or pressure, while this was akin to the coil-globule transition of classical polymers  [23, 24, 25, 26, 27, 28, 29, 30, 31]. In accordance with the reports by Markland et al. [8, 9], increasing the quantumness parameter by compression or cooling promoted the residual diffusion of atoms in the HQDA state. Thus, the NQE is certainly a cause of the polyamorphism and the enhancement of atomic diffusivity in these quantum glasses.
@@
This paper is organized as follows. Section  II describes the CMD simulation method. The results of the isothermal compression simulation are described in Sec. III. At the beginning of this section, we provide the state diagram on the $P$-$T$ plane as a summary of the isothermal compression simulation. And then we show the results of static and dynamic properties supporting the conclusion of this state diagram. In Sec. IV, we present the detailed analysis on the basis of the state diagram. The discussions are given in Sec. V. Finally, we describe the conclusions in Sec. VI. Appendix provides the results of another CMD simulation conducted to examine the occurrence of inverse freezing.
@@ II.1 Isothermal-isobaric CMD
where $v$ is a certain quantity of volume and $\rho_{\rm{c}}$ is the centroid density,
@@ II.2 Check of the Trotter number dependence
First, we checked the Trotter number $N_{\rm{b}}$ dependence of energetic properties. The results of the energetic properties at 0.1 K are shown in Fig. SA-1 in the Supplementary Material. In this figure, we can see a fair convergence of the energetic properties for $N_{\rm{b}}\geq{500}$, so that we adopted $N_{\rm{b}}$ as 500 in all the simulations in the present study. Figure SA-2 in the Supplementary Material shows the Trotter number dependence of energy-temperature relations obtained from the CMD simulation at 1 bar. In this figure, the energetic properties for $T\lesssim 0.1$ K depend fairly on $N_{\rm{b}}$ and the convergence of estimated energy values is not good. For $T\lesssim 0.1$ K, remarkable dropping of $H$, $U$, and $\Phi$ is observed for $500\leq{N_{\rm{b}}}\leq{2000}$, though $K$ converges fairly well. We see that $\Phi$, $U$, and $H$ were evaluated low too much at $T\lesssim{0.1}$ K when we adopted the setting of $N_{\rm{b}}\leq{2000}$. However, there is a symptom that the temperature dependence plot of each energy in the case of $N_{\rm{b}}=3000$ is likely to converge asymptotically to a horizontal value as $T\rightarrow{0}$ K. The plot looks like a slight S-curve in the range of $0.01\leq{T}\leq{0.1}$ K. If we adopted the Trotter number as $N_{\rm{b}}>3000$, more accurate results would be obtained even for $T\lesssim 0.1$ K. Although the low temperature properties at $T<0.1$ K are intriguing, we would rather save this subject for a future work. As for the results shown in this paper, we need to be careful that the thermodynamic properties at $T\lesssim 0.1$ K in this work, where we set $N_{\rm{b}}=500$, contains this degree of inaccuracy.
@@ II.3 Survey of pressure-temperature points by isothermal compression
We surveyed possible states at 357 $P$-$T$ mesh points in the range of $0.08\leq{T}\leq{3.3}$ K and $1\leq{P}\leq{500}$ bar; Figure SA-3 in the Supplementary Material shows these points plotted on the $P$-$T$ plane. The procedure of the CMD simulation, shown schematically in Fig. SA-4 in the Supplementary Material, is as follows:
@@ III.1 State diagram
Figure 1: The state diagram of distinguishable 4He obtained from the isothermal compression by the CMD simulation. The points at 1 bar were obtained from the isobaric cooling from the state at 3.3 K. LQDL (blue): low quantum dispersion liquid; HQDL (red): high quantum dispersion liquid; LQDA (cyan): low quantum dispersion amorphous solid (metastable); HQDA (magenta): high quantum dispersion amorphous solid (metastable). Black solid lines are the experimental phase boundary lines of real 4He. Experimentally, He-I and He-II (gray slant stripe) exist at higher and lower temperatures than $\rm{\lambda}$ transition line (black), respectively, while the crystalline solid phase exists at pressures higher than the melting curve (black). The open symbols denote the states obtained from the isothermal compression in our previous study [21]. The points of “LQDL (Kinugawa et al., 2021)” [21] includes those of supercritical fluid (SCF) state with low quantum dispersion, existing beyond the critical point (red star). The melting line (dashed green line) is presumed to be located at lower pressures than the glass transition line (red line) because of overpressurization. The Kim-Chan phase boundary denotes the locus of the emergence of non-classical rotational inertia which then suggested the existence of supersolid at lower temperatures below this boundary [32]; see the discussion in Sec. V.1.
@@
For visual understanding, in Fig. 2 we show the snapshots of typical instantaneous configurations of LQDL and HQDL at 1 bar. Those for 40 bar are also shown in Fig. SA-5 in the Supplementary Material. In these figures, we find that the LQDA consists of localized necklaces representing low quantum dispersion (LQD) of 4He atoms, while HQDA involves stretched necklaces representing high quantum dispersion (HQD) of atoms. However, the border of the two states is visually ambiguous.
@@ III.2 Thermodynamic properties
In Fig. 4, the slope of the $H$-$T$ plot, $C_{P}$, is larger at lower temperatures. For example, the $C_{P}$ at 1 bar obtained from the curve fitting of the $H$-$T$ relation is shown in Fig. SA-10 in the Supplementary Material. In this figure, we cannot see any indication of such $C_{P}$ maximum as the ${\rm{\lambda}}$-type peak in real 4He and the cusp in the ideal Bose gas [4]. Therefore, there are no evidences for supporting the second-order phase transition between LQDL and HQDL. Consequently, each isobaric transition of HQDL-LQDL, HQDL-LQDA, and HQDA-LQDA is considered as a continuous crossover, not the thermodynamic phase transition.
There is a remarkable drop of $H$ at $T\lesssim{0.1}$ K in Fig. 4, and equivalently, one sees the divergent increase of $C_{P}$ at the lowest temperature edge ($T\lesssim 0.1$ K) in Fig. SA-10 in the Supplementary Material. This is an artifact due to the insufficient $N_{\rm{b}}$ (Sec. II.2). Though the computed values of, e.g., energetic properties, near 0.1 K are not completely accurate, the bulk of this work is hardly affected by the inaccuracy emerging at $T\lesssim 0.1$ K. As described in Sec. II.2, when we adopted the Trotter number $N_{\rm{b}}=3000$, which should yield more accurate results, the $H$-$T$ plot within $0.10.01<T<$ 0.1 K exhibited a symptom of slight S-curve with a subtle inflection point at $\sim 0.02$ K (see Fig. SA-2(e) in the Supplementary Material). Then, the $C_{P}$ may have a maximum at this temperature, and this would possibly imply another state transition. However, we do not draw a conclusion at present; this subject will be resolved in a future study which must demand very heavy computation with enormous $N_{\rm{b}}$.
@@
Figure 5: The kinetic energy vs molar volume. LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). All the experimental data are cited from Fig. 16 of Ref. 50; the measured temperature of normal fluid (He-I) and superfluid (He-II) spans $2.1\leq{T}\leq{2.3}$ K and $0.045\leq{T}\leq{0.075}$ K, respectively [50].

### v3

@@ Abstract
We show that there can exist two liquid states in distinguishable helium-4 (4He) obeying Boltzmann statistics by path integral centroid molecular dynamics (CMD) simulations. This is an indication of quantum liquid polyamorphism induced by nuclear quantum effect. For 0.08-3.3 K and 1-500 bar, we extensively conducted the isothermal-isobaric CMD simulations to explore not only possible states and state diagram but the state characteristics. The distinguishable 4He below 25 bar does not freeze down to 0.1 K even though it includes no Bosonic exchange effect and therefore no Bose condensation. One liquid state, low quantum-dispersion liquid (LQDL), is nearly identical to normal liquid He-I of real 4He. The other is high quantum-dispersion liquid (HQDL) consisting of atoms with longer quantum wavelength. This is another non-superfluid existing below 0.5 K or the temperatures of LQDL. The HQDL is also a low-entropy and fragile liquid to exhibit, unlike conventional liquids, rather gas-like relaxation of velocity autocorrelation function, while there the atoms diffuse without noticeable contribution from quantum tunneling. The LQDL-HQDL transition is not a thermodynamic phase transition but a continuous crossover accompanied by the change of the expansion factor of quantum wavelength. Freezing of HQDL into the low quantum-dispersion amorphous solid occurs by heating from 0.2 to 0.3 K at 40-50 bar, while this $P$-$T$ condition coincides with the Kim-Chan normal-supersolid phase boundary of real 4He. The obtained state diagram was compared with that of the confined subnano-scale 4He systems where Bosonic correlation is considerably suppressed.
helium-4, polyamorphism, path integral, nuclear quantum effect, molecular dynamics, inverse freezing, inverse melting, quantum liquid
@@ I INTRODUCTION
Is it a true postulation that every substance should freeze into its solid phase as it is cooled down toward absolute zero temperature? While classical systems inevitably freeze by losing kinetic energy proportional to temperature, anti-freezing of quantum systems is benefited by promotion of atomic diffusivity due to increasing quantum atomic fluctuation down to zero temperature. Indeed, we know an exceptionally non-freezing quantum substance, helium-4 (4He), which retains the superfluid (He-II) phase which never freezes down to zero temperature under atmospheric pressure. The cause of superfluidity includes: (i) the nuclear quantum effect (NQE), i.e., the atomic wave nature parametrized as the de Broglie thermal wavelength $\lambda_{\rm{dB}}:=\hbar\sqrt{2\pi/mk_{\rm{B}}T}$ ($m$ is the atomic mass and $k_{\rm{B}}$ is the Boltzmann constant), (ii) the weakness of the van der Waals interatomic forces, and (iii) the atomic indistinguishableness due to the Bosonic permutation [1]. Primitively, the $\rm{\lambda}$-type isobaric heat capacity $C_{P}$ at the transition between normal liquid (He-I) and He-II of 4He is explained in terms of the Bose-Einstein condensation (BEC) of ideal Bose gas [2, 3, 4]. This leads to a conventional understanding that the Bose statistics is the key to the $\lambda$ transition and causes the BEC through which non-freezing of this quantum substance is achieved. Then, if there were solely the NQE and no Bosonic exchange were imposed on atoms, could such a distinguishable 4He model avoid freezing to retain the normal liquid phase when cooled? According to Feynman [1], it was argued that there was no difference between the ground state wave function of an ensemble of Bosons and of distinguishable particles [5]. This indicates that distinguishable 4He at atmospheric pressure is liquid at $T=0$ as is the Bosonic system. As discussed [5], in the textbooks [1, 6], the non-freezing property of 4He was explained by considering only the two factors (i) the NQE of individual atoms and (ii) the weak interatomic forces [6]. However, Boninsegni et al. reported that the third factor, (iii) the Bosonic exchange among $N$ atoms, was indispensable to retaining the non-freezing property by being in superfluid state; the crystal of distinguishable 4He melted by introducing the Bosonic correlation [5]. Therefore, the state of distinguishable 4He would not always be the same as that of the Bosonic counterpart in general. However, for a wider range of pressure and non-zero low temperature, it has not yet been completely explored what kind of states emerges by the sole NQE in distinguishable 4He.
@@
The non-freezing tendency of a quantum substance might be reflected to the ease of residual diffusion [7] of atoms in its quenched glassy solid. In this connection, Markland et al. conducted the discretized path integral simulations based on the quantum-classical isomorphism, where quantum atoms are expressed as classical ring polymers (necklaces), to show the reentrant behavior of self-diffusion coefficient $D$ in the glassy solid of a binary Lennard-Jones model. In their study, increasing the quantumness parameter $\Lambda^{*}:=\hbar/\sqrt{mk_{\rm{B}}T\sigma^{2}}$ ($\sigma$ is the closest interatomic distance) caused the transition from a glass state involving trapped configurations of atomic necklaces (called trapped regime) to another state with stretched configurations of atoms (tunneling regime) [8, 9]. As $\Lambda^{*}$ increased, $D$ once decreased in the former regime but it turned to increase in the latter [8, 9]. This also suggested a tendency toward inverse melting (or melting by cooling) of this quantum system [9]. In fact, such an atomic tunneling between local potential minima explains the linear temperature dependence of heat capacity in low-temperature glasses [10, 11].
@@
The existence of more than one states in liquids or glasses is called polyamorphism [12, 13], which is, for example, observed for classical fragile liquids [14] such as water [15, 16] and phosphorous [17]. The existence of two liquid states, He-I and He-II, in 4He is certainly an example of liquid polyamorphism. However, the polyamorphism of the other quantum systems has rarely been discovered so far [18, 19, 20]. Recently, our path integral centroid molcular dynamics (CMD) simulations showed a quantum polyamorphism of compressed distinguishable 4He glass above 3 K [21]. In good accordance with the two regimes reported by Markland et al. [8, 9], compressed distinguishable 4He also involved two glass states [21]. The two states were distinguished by the difference in atomic quantum wavelength $\lambda_{\rm{quantum}}$ which is equivalent to two times radius of gyration of atomic necklaces [21]. Then we call the two states, depending on $\lambda_{\rm{quantum}}$, the low quantum dispersion amorphous solid (LQDA) and the high quantum dispersion solid (HQDA) [21]  [^1] Here we should mention the abbreviation of these names. In Ref. 21, we abbreviated them as “LDA” and “HDA”, respectively. However, since these may be confused with the low and high density amorphous state of classical systems (e.g., water [15, 16]), we presently call them LQDA and HQDA instead. The LQDA and HQDA are equivalent to the trapped and the tunneling regimes in the quantum Lennard-Jones model [8, 9], respectively. The LQDA-HQDA transition was reversibly caused by varying temperature or pressure, while this was akin to the coil-globule transition of classical polymers [23, 24, 25, 26, 27, 28, 29, 30, 31]. In accordance with the reports by Markland et al. [8, 9], increasing the quantumness parameter by compression or cooling promoted the residual diffusion of atoms in the HQDA state. Thus, the NQE is certainly a cause of the polyamorphism and the enhancement of atomic diffusivity in these quantum glasses.
@@
This paper is organized as follows. Section  II describes the CMD simulation method. The results of the isothermal compression simulation are described in Sec. III. At the beginning of this section, we provide the state diagram on the $P$-$T$ plane as a summary of the isothermal compression simulation. And then we show the results of static and dynamic properties supporting the conclusion of this state diagram. In Sec. IV, we present the detailed analysis on the basis of the state diagram. The discussions are given in Sec. V. Finally, we describe the conclusions in Sec. VI. Appendix provides the results of another CMD simulation conducted to examine the occurrence of inverse freezing, and also gives comparisons of the state diagram with that of the real confined 4He systems and reported supersolid.
@@ II.1 Isothermal-isobaric CMD
where $v$ is a certain quantity of volume [44] and $\rho_{\rm{c}}$ is the centroid density,
@@ II.2 Check of the Trotter number dependence
First, we checked the Trotter number $N_{\rm{b}}$ dependence of energetic properties. The results of the energetic properties at 0.1 K are shown in Fig. SA-1 in the Supplementary Material. In this figure, we can see a fair convergence of the energetic properties for $N_{\rm{b}}\geq{500}$, so that we adopted $N_{\rm{b}}$ as 500 in all the simulations in the present study. Figure SA-2 in the Supplementary Material shows the Trotter number dependence of energy-temperature relations obtained from the CMD simulation at 1 bar. In this figure, the energetic properties for $T\lesssim 0.1$ K depend fairly on $N_{\rm{b}}$ and the convergence of estimated energy values is not good. For $T\lesssim 0.1$ K, remarkable dropping of $H$, $U$, and $\Phi$ is observed for $500\leq{N_{\rm{b}}}\leq{2000}$, though $K$ converges fairly well. We see that $\Phi$, $U$, and $H$ were evaluated low too much at $T\lesssim{0.1}$ K when we adopted the setting of $N_{\rm{b}}\leq{2000}$. However, there is a symptom that the temperature dependence plot of each energy in the case of $N_{\rm{b}}=3000$ is likely to converge asymptotically to a horizontal value as $T\rightarrow{0}$ K. The plot looks like a slight S-curve in the range of $0.01\leq{T}\leq{0.1}$ K. If we adopted the Trotter number as $N_{\rm{b}}>3000$, more accurate results would be obtained even for $T\lesssim 0.1$ K. Although the low temperature properties at $T<0.1$ K are intriguing, we would rather save this subject for a future work. As for the results shown in this paper, we need to be careful that the thermodynamic properties at $T\lesssim 0.1$ K in this work, where we set $N_{\rm{b}}=500$, contains this degree of inaccuracy. The $N_{\rm{b}}$-dependence of other properties is also shown in Figs. SA-3 and SA-4 in the Supplementary Material; this issue is discussed in Secs. III.3 and III.4.
@@ II.3 Survey of pressure-temperature points by isothermal compression
We surveyed possible states at 357 $P$-$T$ mesh points in the range of $0.08\leq{T}\leq{3.3}$ K and $1\leq{P}\leq{500}$ bar; Figure SA-5 in the Supplementary Material shows these points plotted on the $P$-$T$ plane. The procedure of the CMD simulation, shown schematically in Fig. SA-6 in the Supplementary Material, is as follows:
@@ III.1 State diagram
Figure 1: The state diagram of distinguishable 4He obtained from the isothermal compression by the CMD simulation. The points at 1 bar were obtained from the isobaric cooling from the state at 3.3 K. LQDL (blue): low quantum dispersion liquid; HQDL (red): high quantum dispersion liquid; LQDA (cyan): low quantum dispersion amorphous solid (metastable); HQDA (magenta): high quantum dispersion amorphous solid (metastable). Black solid lines are the experimental phase boundary lines of real 4He. Experimentally, He-I and He-II (gray slant stripe) exist at higher and lower temperatures than $\rm{\lambda}$ transition line (black), respectively, while the crystalline solid phase exists at pressures higher than the melting curve (black). The open symbols denote the states obtained from the isothermal compression in our previous study [21]. The points of “LQDL (Kinugawa et al., 2021)” [21] includes those of supercritical fluid (SCF) state with low quantum dispersion, existing beyond the critical point (red star). The melting line (dashed green line) is presumed to be located at lower pressures than the glass transition line (red line) because of overpressurization. The Kim-Chan phase boundary denotes the locus of the emergence of non-classical rotational inertia which then suggested the existence of supersolid at lower temperatures below this boundary [32]; see Appendix B.
@@
For visual understanding, in Fig. 2 we show the snapshots of typical instantaneous configurations of LQDL and HQDL at 1 bar. Those for 40 bar are also shown in Fig. SA-7 in the Supplementary Material. In these figures, we find that the LQDA consists of localized necklaces representing low quantum dispersion (LQD) of 4He atoms, while HQDA involves stretched necklaces representing high quantum dispersion (HQD) of atoms. However, the border of the two states is visually ambiguous.
@@ III.2 Thermodynamic properties
In Fig. 4, the slope of the $H$-$T$ plot, $C_{P}$, is larger at lower temperatures. For example, the $C_{P}$ at 1 bar obtained from the curve fitting of the $H$-$T$ relation is shown in Fig. SA-12 in the Supplementary Material. In this figure, we cannot see any indication of such $C_{P}$ maximum as the ${\rm{\lambda}}$-type peak in real 4He and the cusp in the ideal Bose gas [4]. Therefore, there are no evidences for supporting the second-order phase transition between LQDL and HQDL. Consequently, each isobaric transition of HQDL-LQDL, HQDL-LQDA, and HQDA-LQDA is considered as a continuous crossover, not the thermodynamic phase transition.
There is a remarkable drop of $H$ at $T\lesssim{0.1}$ K in Fig. 4, and equivalently, one sees the divergent increase of $C_{P}$ at the lowest temperature edge ($T\lesssim 0.1$ K) in Fig. SA-12 in the Supplementary Material. This is an artifact due to the insufficient $N_{\rm{b}}$ (Sec. II.2). Though the computed values of, e.g., energetic properties, near 0.1 K are not completely accurate, the bulk of this work is hardly affected by the inaccuracy emerging at $T\lesssim 0.1$ K. As described in Sec. II.2, when we adopted the Trotter number $N_{\rm{b}}=3000$, which should yield more accurate results, the $H$-$T$ plot within $0.10.01<T<$ 0.1 K exhibited a symptom of slight S-curve with a subtle inflection point at $\sim 0.02$ K (see Fig. SA-2(e) in the Supplementary Material). Then, the $C_{P}$ may have a maximum at this temperature, and this would possibly imply another state transition. However, we do not draw a conclusion at present; this subject will be resolved in a future study which must demand very heavy computation with enormous $N_{\rm{b}}$.
@@
Figure 5: The kinetic energy vs molar volume. LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). All the experimental data are cited from Fig. 16 of Ref. 51; the measured temperature of normal fluid (He-I) and superfluid (He-II) spans $2.1\leq{T}\leq{2.3}$ K and $0.045\leq{T}\leq{0.075}$ K, respectively [51].